# Evolving Generalizable Parallel Algorithm Portfolios via Domain-Agnostic Instance Generation

### v3

@@ Abstract
Index Terms:
Algorithm configuration, parallel algorithm portfolios, automatic algorithm design, co-evolutionary algorithm, binary optimization problem
@@ I Introduction
In recent decades, search-based methods such as Evolutionary Algorithms (EAs) have become the mainstream approach for solving NP-hard optimization problems [1, 2, 3, 4]. Most, if not all, of these methods involve a set of free parameters that would affect their search behavior. While theoretical analyzes for many search-based methods have offered worst or average bounds on their performance, their actual performance in practice is highly sensitive to the settings of parameters, i.e., algorithm configurations. However, finding the optimal algorithm configurations requires both domain-specific knowledge and algorithmic expertise, which cannot be done manually with ease. In response to this, significant efforts have been made to automate the tuning procedure, commonly referred to as automatic parameter tuning [5] and automatic algorithm configuration (AAC) [6, 7, 8]. Typically, AAC follows a high-level generate-and-evaluate process, where different configurations are iteratively generated and evaluated on a given set of problem instances, i.e., the training set. Upon termination, the best-performing configuration is returned. Since an algorithm configuration fully instantiates a parameterized algorithm, for brevity, throughout this article we will use the term “configuration” to refer to the resultant algorithm.
@@
Building upon AAC, a series of works have been conducted to identify a set of configurations instead of a single configuration to form a parallel algorithm portfolio (PAP), referred to as automatic construction of PAPs [9, 10, 11, 12]. Each configuration in the PAP is called a member algorithm. When applied to a problem instance, all member algorithms in the PAP run independently in parallel to obtain multiple solutions, from which the best solution is returned. Although a PAP consumes more computational resources than a single algorithm, it can achieve superior overall performance than any single algorithm through the complementary strengths of its member algorithms [13, 14]. That is, different member algorithms of the PAP excel at solving different types of problem instances. Moreover, PAPs employ simple parallel solution strategies, making them well-suited to exploit modern computing facilities (e.g., multi-core CPUs). Nowadays, such capability has become increasingly crucial for tackling computationally challenging problems [15], given the tremendous advancement of parallel computing architectures in the past decade [16].
@@
Generalization is the core objective in the automatic construction of PAPs [11]. It requires that the constructed PAP performs well not only on instances within the training set but also on unseen instances from the same problem class. Given a training set that sufficiently represents the problem class, existing approaches have proven highly effective in constructing PAPs with good generalization [9, 17, 18, 10]. However, in real-world applications, one is very likely to encounter the few-shots scenarios, where the collected training instances are limited and biased (e.g., sampled from a specific distribution) and thus fail to sufficiently represent the target problem class [19]. To address this challenge, recent studies have explored integrating instance generation into the construction process [11, 12, 20]. One representative approach is the co-evolution of parameterized search (CEPS) [11] that maintains two competing populations during evolution: a configuration population (PAP) and a problem instance population. The evolution of the instance population aims to identify challenging instances that exploit weaknesses in the current PAP, while the evolution of the PAP improves its generalization by identifying configurations that better handle these instances.
@@
Although CEPS has shown promising performance in few-shot scenarios, it has one major limitation. That is, CEPS fundamentally relies on domain-specific instance generators, which significantly limits its generality and practical applicability. As shown in Fig. 1, when applying CEPS to a specific problem class, one needs to provide instance generators tailored to the problem class to generate problem instances during the evolution of the instance population. As demonstrated in [11], when applying CEPS to the traveling salesman problem (TSP), Tang et al. developed a TSP-specific instance mutation operator based on 2D coordinate representations of TSP instances. However, designing such generators requires comprehensive knowledge of the problem class, including suitable instance representations and corresponding variation mechanisms. This becomes particularly challenging for newly emerging optimization problems and black-box problems where such knowledge is not readily available. For example, consider the compiler argument optimization problem [21], which aims to find the optimal compiler argument settings to minimize the size of the binary file generated from compiling a given program source code. For such a problem, the instance generator should be able to create new yet valid program source code – a task that demands deep understanding of program syntax and semantics.
@@
This work proposes a novel PAP construction approach called domain-agnostic co-evolution of parameterized search (DACE) that eliminates the need for domain-specific instance generators. DACE is applicable to binary optimization problems where decision variables take values of 0 or 1. Such problems are widespread in practice, e.g., appearing in combinatorial optimization [^1] Theoretically, any discrete optimization problem can be reformulated as a binary optimization problem., discrete facility location, and certain graph problems (e.g., max cut and max coverage).
@@
    The proposed domain-agnostic instance representation and generation approach serves as a general data augmentation technique, which not only benefits PAP construction but also opens up new possibilities for training broad classes of optimizers, i.e., learning to optimize (L2O) [22].
@@ II-A Notations and Problem Description
Assume a PAP is to be built for a problem class, for which we denote a problem instance as $s$ and the set of all possible instances as $\Omega$. Given a parameterized optimization algorithm with its configuration space $\Theta$, each $\theta\in\Theta$ is an algorithm configuration that fully instantiates the algorithm. Let $P=\left\{\theta_{1},\theta_{2},\cdots,\theta_{K}\right\}$ denote a PAP that consists of $K$ configurations as its member algorithms. For any problem instance $s\in\Omega$ and configuration $\theta\in\Theta$, let $f\left(\theta,s\right)$ denote the performance of $\theta$ on $s$. In practice, this indicator can measure various aspects of performance such as solution quality [7], computational efficiency [23], or even be stated in a multi-objective form [24]. Without loss of generality, we assume larger values indicate better performance. The performance of $P$ on instance $s$ is the best performance achieved among its member algorithms $\theta_{1},\theta_{2},\cdots,\theta_{K}$ on $s$:
@@ II-B Existing Approaches for Constructing Generalizable PAPs
Existing approaches such as GLOBAL [9], PARHYDRA [18], CLUSTERING [17], and PCIT [10] have proven effective in such scenarios. However, when the training instances are rather limited, i.e., few-shot scenarios, optimizing performance solely on the training set can lead to the overtuning phenomenon [23], similar to the overfitting in machine learning. That is, the constructed PAP performs well on the training set but its test (generalization) performance is arbitrarily bad.
Leveraging synthetic instances during the construction process has been shown to be effective in addressing the above challenge [11, 12, 20]. The representative approach is CEPS [11] that employs a co-evolutionary framework [25] where a configuration population $P$ and an instance population $T$ compete with each other. Specifically, each round of its co-evolution has two steps:
@@
As shown in [11], the two-step process effectively maximizes the lower bound, i.e., a tractable surrogate, of PAP’s generalization performance as defined in Eq. (3). While CEPS has shown promising results in few-shot scenarios, applying it to a specific problem class requires one to provide domain-specific instance generators, particularly mutation operators, for evolving instances in the first step.
@@ II-C Existing Problem Instance Generation Methods
Designing effective instance generators is crucial for the few-shot construction of generalizable PAPs, yet it is often difficult, as discussed earlier. Domain-specific generators developed for particular problems, such as TSP [11, 26], vehicle routing problem [11], and job shop scheduling problem [27], can be utilized but are limited only to those specific problem domains. In deep learning, instance generation methods have also been developed to address few-shot scenarios in vision  [28] and natural language domains [29]. However, due to the fundamental differences between these domains (images and text) and optimization problems, these approaches cannot be applied to the latter. Furthermore, for emerging and black-box optimization problems where even experts lack sufficient domain knowledge, designing domain-specific instance generators becomes even more challenging. The following sections will present DACE that eliminates the need for domain-specific instance generators.
@@ III Domain-Agnostic Problem Instance Representation and Generation
The key innovation of DACE lies in its domain-agnostic approach to representing and generating problem instances. At its core, DACE employs neural networks (NNs) as a universal representation for problem instances, dubbed neural instance representation (NIR). Given a small set $T$ of training instances, DACE first converts these instances into NIRs and then uses these NIRs as a basis for generating new instances, which are also represented as NIRs. Since different problem instances essentially differ in their underlying objective functions, specifically, in how they map solutions to objective values. Therefore, an NIR represents a problem instance by approximating its objective function. This way, the training process of an NIR only requires pairs of solutions and their corresponding objective values, without requiring any domain-specific knowledge such as analytic form of the objective function. This means training instances in $T$ can be provided as black boxes that simply return solution evaluations. Furthermore, by varying the weight parameters within the NIR, different objective functions corresponding to different problem instances are obtained.
@@
Technically, NIRs share similarities with surrogate models, a technique widely used in EAs to reduce the number of expensive fitness evaluations [30, 31, 32, 33]. However, it is important to note that NIRs serve a fundamentally different purpose here. Rather than reducing number of fitness evaluations, NIRs serve as the basis for instance generation, where all generated instances in DACE are represented as NIRs instead of their actual forms. To achieve this goal, two technical challenges must be addressed. First, effective mutation operators need to be developed that can generate meaningful instances represented as NIRs. Due to the powerful representation capabilities of NNs, particularly deep NNs, random mutations of NIR parameters could produce arbitrary objective functions that do not correspond to valid actual problem instances. In other words, the generated instances should relate to the training instances and belong to the same problem class. This requires the extraction and utilization of domain-invariant features from the training instances to properly constrain the instance generation process. The second challenge arises from the discrete nature of binary optimization problems. In these problems, small changes in discrete inputs (such as flipping a few bits) can result in dramatic changes in objective values. This makes NN learning particularly challenging, as NNs are typically suited for fitting smooth, continuous functions [34]
@@
To address these challenges, we propose a decoupled design of the structure of NIR based on variational autoencoders (VAE) [35] and hypernetworks [36]. This structure decouples domain-invariant features from instance-specific features, and also decouples solution encoding from objective function approximation. Below we detail the NIR structure, its training method, and the NIR-based instance mutation operator.
@@ III-A Structure of the NIR
Let $\boldsymbol{w}_{S}$ denote the weight parameters of $F_{S}$, the concatenation (denoted by $\oplus$)⊕ ) of $\boldsymbol{\mu}_{z}$ and $\boldsymbol{\sigma}_{z}$ is fed into $F_{S}$, which predicts the score (objective value) $y^{\prime}$:
@@
This means the weight parameters $\boldsymbol{w}_{S}$ of the scorer $F_{S}$ is generated by the hypernetwork $F_{H}$ through its output conditioned on the input embedding $\boldsymbol{e}$.
@@ III-B Training NIRs
Here, $\boldsymbol{\mu}_{z},\boldsymbol{\sigma}_{z}$ are obtained through Eq. (5), and $\boldsymbol{x}^{\prime},y^{\prime}$ are obtained through Eq. (6) and Eq. (7), respectively. The loss function consists of three terms. The first term is a reconstruction loss between the input solution $\boldsymbol{x}$ and its reconstructed counterpart $\boldsymbol{x}^{\prime}$, promoting the encoder to capture the structure of $\boldsymbol{x}$. The second term is a prediction loss for objective values, encouraging the scorer to be accurate. The third term is the KL-divergence between the learned latent distribution and the standard Gaussian distribution, which serves as a standard regularization term commonly used in VAEs [35] to ensure a smooth latent space. All parameters are randomly initialized and jointly optimized using stochastic gradient descent.
Input: Problem instance represented as NIR $m$, PAP $P$, black-box continuous optimizer $OPT$.
Output: Mutated instance represented as NIR $m^{\prime}$.
1 Initialize $OPT$ with instance embedding $\boldsymbol{e}$ of $m$;
2 $\boldsymbol{e}^{\prime}\leftarrow$ minimize Eq. (10) using $OPT$;
4 return $m^{\prime}\textrm{ specified by }\boldsymbol{e}^{\prime}$;
Algorithm 1 NIR-based Instance Mutation Operator
Input: Training set $T$; number of member algorithms $K$; maximum number of configuration mining iterations $n$; maximum round number $MaxRound$.

### v4

# Evolving Generalizable Parallel Algorithm Portfolios for Binary Optimization Problems via Domain-Agnostic Instance Generation

@@ I Introduction
In recent decades, search-based methods such as Evolutionary Algorithms (EAs) have become the mainstream approach for solving NP-hard optimization problems [wang2022genetic, zhou2021survey, li2017seeking, beke2024routing]. Most, if not all, of these methods involve a set of free parameters that would affect their search behavior. While theoretical analyzes for many search-based methods have offered worst or average bounds on their performance, their actual performance in practice is highly sensitive to the settings of parameters, i.e., algorithm configurations. However, finding the optimal algorithm configurations requires both domain-specific knowledge and algorithmic expertise, which cannot be done manually with ease. In response to this, significant efforts have been made to automate the tuning procedure, commonly referred to as automatic parameter tuning [huang_survey_2020] and automatic algorithm configuration (AAC) [ansotegui_model-based_2015, lopez-ibanez_irace_2016, hutter_sequential_2011]. Typically, AAC follows a high-level generate-and-evaluate process, where different configurations are iteratively generated and evaluated on a given set of problem instances, i.e., the training set. Upon termination, the best-performing configuration is returned. Since an algorithm configuration fully instantiates a parameterized algorithm, for brevity, throughout this article we will use the term “configuration” to refer to the resultant algorithm.
@@
Building upon AAC, a series of works have been conducted to identify a set of configurations instead of a single configuration to form a parallel algorithm portfolio (PAP), referred to as automatic construction of PAPs [lindauer_automatic_2017, liu_automatic_2019, tang_few-shots_2021, liu_generative_2022]. Each configuration in the PAP is called a member algorithm. When applied to a problem instance, all member algorithms in the PAP run independently in parallel to obtain multiple solutions, from which the best solution is returned. Although a PAP consumes more computational resources than a single algorithm, it can achieve superior overall performance than any single algorithm through the complementary strengths of its member algorithms [huberman1997economics, gomes_algorithm_2001]. That is, different member algorithms of the PAP excel at solving different types of problem instances. Moreover, PAPs employ simple parallel solution strategies, making them well-suited to exploit modern computing facilities (e.g., multi-core CPUs). Nowadays, such capability has become increasingly crucial for tackling computationally challenging problems [sutton2019bitter], given the tremendous advancement of parallel computing architectures in the past decade [hockney2019parallel].
@@
Generalization is the core objective in the automatic construction of PAPs [tang_few-shots_2021]. It requires that the constructed PAP performs well not only on instances within the training set but also on unseen instances from the same problem class. Given a training set that sufficiently represents the problem class, existing approaches have proven highly effective in constructing PAPs with good generalization [lindauer_automatic_2017, kadioglu_isac_2010, xu_hydra_2010, liu_automatic_2019]. However, in real-world applications, one is very likely to encounter the few-shots scenarios, where the collected training instances are limited and biased (e.g., sampled from a specific distribution) and thus fail to sufficiently represent the target problem class [smith-miles_generating_2015]. To address this challenge, recent studies have explored integrating instance generation into the construction process [tang_few-shots_2021, liu_generative_2022, wang_asp_2024]. One representative approach is the co-evolution of parameterized search (CEPS) [tang_few-shots_2021] that maintains two competing populations during evolution: a configuration population (PAP) and a problem instance population. The evolution of the instance population aims to identify challenging instances that exploit weaknesses in the current PAP, while the evolution of the PAP improves its generalization by identifying configurations that better handle these instances.
@@
Although CEPS has shown promising performance in few-shot scenarios, it has one major limitation. That is, CEPS fundamentally relies on domain-specific instance generators, which significantly limits its generality and practical applicability. As shown in Fig. 1, when applying CEPS to a specific problem class, one needs to provide instance generators tailored to the problem class to generate problem instances during the evolution of the instance population. As demonstrated in [tang_few-shots_2021], when applying CEPS to the traveling salesman problem (TSP), Tang et al. developed a TSP-specific instance mutation operator based on 2D coordinate representations of TSP instances. However, designing such generators requires comprehensive knowledge of the problem class, including suitable instance representations and corresponding variation mechanisms. This becomes particularly challenging for newly emerging optimization problems and black-box problems where such knowledge is not readily available. For example, consider the compiler argument optimization problem [jiang_smartest_2022], which aims to find the optimal compiler argument settings to minimize the size of the binary file generated from compiling a given program source code. For such a problem, the instance generator should be able to create new yet valid program source code – a task that demands deep understanding of program syntax and semantics.
@@
This work proposes a novel PAP construction approach called domain-agnostic co-evolution of parameterized search (DACE). The term “domain-agnostic” refers specifically to the problem domain. That is, unlike existing methods requiring domain-specific instance generators built upon a problem instance’s structural information (e.g., the TSP instance generator as described above), DACE has a general-purpose generator that does not need such information. This allows DACE to be applicable to diverse binary optimization problem classes where variables take 0/1 values. Such problems are widespread in practice, e.g., appearing in combinatorial optimization1, discrete facility location, and certain graph problems (e.g., max cut and max coverage).
@@
    The proposed domain-agnostic instance representation and generation approach serves as a general data augmentation technique, which not only benefits PAP construction but also opens up new possibilities for training broad classes of optimizers, i.e., learning to optimize (L2O) [tang2024learn].
@@ II-A Notations and Problem Description
Assume a PAP is to be built for a problem class, for which we denote a problem instance as $s$ and the set of all possible instances as $\Omega$. Given a parameterized optimization algorithm with its configuration space $\Theta$, each $\theta\in\Theta$ is an algorithm configuration that fully instantiates the algorithm. Let $P=\left\{\theta_{1},\theta_{2},\cdots,\theta_{K}\right\}$ denote a PAP that consists of $K$ configurations as its member algorithms. For any problem instance $s\in\Omega$ and configuration $\theta\in\Theta$, let $f\left(\theta,s\right)$ denote the performance of $\theta$ on $s$. In practice, this indicator can measure various aspects of performance such as solution quality [lopez-ibanez_irace_2016], computational efficiency [liu_performance_2020], or even be stated in a multi-objective form [bolt2016moparamils]. Without loss of generality, we assume larger values indicate better performance. The performance of $P$ on instance $s$ is the best performance achieved among its member algorithms $\theta_{1},\theta_{2},\cdots,\theta_{K}$ on $s$:
@@ II-B Existing Approaches for Constructing Generalizable PAPs
Existing approaches such as GLOBAL [lindauer_automatic_2017], PARHYDRA [xu_hydra_2010], CLUSTERING [kadioglu_isac_2010], and PCIT [liu_automatic_2019] have proven effective in such scenarios. However, when the training instances are rather limited, i.e., few-shot scenarios, optimizing performance solely on the training set can lead to the overtuning phenomenon [liu_performance_2020], similar to the overfitting in machine learning. That is, the constructed PAP performs well on the training set but its test (generalization) performance is arbitrarily bad.
Leveraging synthetic instances during the construction process has been shown to be effective in addressing the above challenge [tang_few-shots_2021, liu_generative_2022, wang_asp_2024]. The representative approach is CEPS [tang_few-shots_2021] that employs a co-evolutionary framework [ma2018survey] where a configuration population $P$ and an instance population $T$ compete with each other. Specifically, each round of its co-evolution has two steps:
@@
As shown in [tang_few-shots_2021], the two-step process effectively maximizes the lower bound, i.e., a tractable surrogate, of PAP’s generalization performance as defined in Eq. (3). While CEPS has shown promising results in few-shot scenarios, applying it to a specific problem class requires one to provide domain-specific instance generators, particularly mutation operators, for evolving instances in the first step.
@@ II-C Existing Problem Instance Generation Methods
Designing effective instance generators is crucial for the few-shot construction of generalizable PAPs, yet it is often difficult, as discussed earlier. Domain-specific generators developed for particular problems, such as TSP [tang_few-shots_2021, Hemert06], vehicle routing problem [tang_few-shots_2021], and job shop scheduling problem [branke2011evolutionary], can be utilized but are limited only to those specific problem domains. In deep learning, instance generation methods have also been developed to address few-shot scenarios in vision  [zhang2022fewshot] and natural language domains [zhou2022fipda]. However, due to the fundamental differences between these domains (images and text) and optimization problems, these approaches cannot be applied to the latter. Furthermore, for emerging and black-box optimization problems where even experts lack sufficient domain knowledge, designing domain-specific instance generators becomes even more challenging. The following sections will present DACE that eliminates the need for domain-specific instance generators.
@@ III Domain-Agnostic Problem Instance Representation and Generation
The key innovation of DACE lies in its domain-agnostic approach to representing and generating problem instances. At its core, DACE employs neural networks (NNs) as a universal representation for problem instances, dubbed neural instance representation (NIR). Given a small set $T$ of training instances, DACE first converts these instances into NIRs and then uses these NIRs as a basis for generating new instances, which are also represented as NIRs. Since different problem instances essentially differ in their underlying objective functions, specifically, in how they map solutions to objective values. Therefore, an NIR represents a problem instance by approximating its objective function. This way, the training process of an NIR only requires pairs of solutions and their corresponding objective values, without requiring any domain-specific knowledge such as analytic form of the objective function. Furthermore, by varying the weight parameters within the NIR, different objective functions corresponding to different problem instances are obtained.
As a result, NIR maintains generality across various binary optimization problem classes, where training instances in $T$ can be treated as black boxes that simply return solution evaluations. It is worth noting that the term “generality” differs from “generalization”. Specifically, PAP’s generalization indicates its ability to perform well not only on training instances, but also on unseen instances from the same problem class, whereas NIR’s generality means its broad applicability to various binary optimization problem classes.
@@
Technically, NIRs share similarities with surrogate models, a technique widely used in EAs to reduce the number of expensive fitness evaluations [Jin11, zhang2021surrogate, nguyen2024constrained, lin2022ensemble]. However, it is important to note that NIRs serve a fundamentally different purpose here. Rather than reducing number of fitness evaluations, NIRs serve as the basis for instance generation, where all generated instances in DACE are represented as NIRs instead of their actual forms. To achieve this goal, two technical challenges must be addressed. First, effective mutation operators need to be developed that can generate meaningful instances represented as NIRs. Due to the powerful representation capabilities of NNs, particularly deep NNs, random mutations of NIR parameters could produce arbitrary objective functions that do not correspond to valid actual problem instances. In other words, the generated instances should relate to the training instances and belong to the same problem class. This requires the extraction and utilization of domain-invariant features from the training instances to properly constrain the instance generation process. The second challenge arises from the discrete nature of binary optimization problems. In these problems, small changes in discrete inputs (such as flipping a few bits) can result in dramatic changes in objective values. This makes NN learning particularly challenging, as NNs are typically suited for fitting smooth, continuous functions [ferrari2005smooth].
@@
To address these challenges, we propose a decoupled design of the structure of NIR based on variational autoencoders (VAE) [kingma_auto-encoding_2014] and hypernetworks [chauhan_brief_2024]. This structure decouples domain-invariant features from instance-specific features, and also decouples solution encoding from objective function approximation. Below we detail the NIR structure, its training method, and the NIR-based instance mutation operator.
@@ III-A Structure of the NIR
Let $\boldsymbol{w}_{S}$ denote the weight parameters of $F_{S}$, the concatenation (denoted by $\oplus)$ of $\boldsymbol{\mu}_{z}$ and $\boldsymbol{\sigma}_{z}$ is fed into $F_{S}$, which predicts the score (objective value) $y^{\prime}$:
@@
This means the weight parameters $\boldsymbol{w}_{S}$ of the scorer $F_{S}$ is generated by the hypernetwork $F_{H}$ through its output conditioned on the input embedding $\boldsymbol{e}$. The instance embedding $\boldsymbol{e}$ is a trainable vector optimized to encode instance-specific information. It is randomly initialized and then trained with (solution, objective value) pairs sampled from the corresponding training problem instance, as detailed in Section III-B.
@@
Input: Problem instance represented as NIR $m$, PAP $P$, black-box continuous optimizer OPTOPT.
Output: Mutated instance represented as NIR $m^{\prime}$.
1 Initialize OPTOPT with instance embedding $\boldsymbol{e}$ of $m$;
2 $\boldsymbol{e}^{\prime}\leftarrow$ minimize Eq. (10) using OPTOPT;
4 return $m^{\prime}\textrm{ specified by }\boldsymbol{e}^{\prime}$;
Algorithm 1 NIR-based Instance Mutation Operator
@@ III-B Training NIRs
Here, $\boldsymbol{\mu}_{z},\boldsymbol{\sigma}_{z}$ are obtained through Eq. (5), and $\boldsymbol{x}^{\prime},y^{\prime}$ are obtained through Eq. (6) and Eq. (7), respectively. The loss function consists of three terms. The first term is a reconstruction loss between the input solution $\boldsymbol{x}$ and its reconstructed counterpart $\boldsymbol{x}^{\prime}$, promoting the encoder to capture the structure of $\boldsymbol{x}$. The second term is a prediction loss for objective values, encouraging the scorer to be accurate. It should be noted that the actual objective value $y$ has been normalized to the range $\left[0,1\right]$ to prevent the influence of varying value scales across different problem instances on the training process. The third term is the KL-divergence between the learned latent distribution and the standard Gaussian distribution, which serves as a standard regularization term commonly used in VAEs [kingma_auto-encoding_2014] to ensure a smooth latent space. All parameters are randomly initialized and jointly optimized using stochastic gradient descent.
During the training process of NIR, constraint information from problem instances is not incorporated, meaning NIR does not directly handle constraints. This is because constraints in optimization problems are typically problem-dependent, and encoding them into the model’s representation would compromise the generality of NIR. To address constrained optimization problems, constraint-handling techniques are applied to convert infeasible solutions into feasible ones before they are fed into NIR. More details can be found in Section V-A.
Input: Training set $T$; number of member algorithms $K$; maximum number of configuration mining iterations $n$; maximum round number MaxRoundMaxRound.